# Feedback-Aware Monte Carlo Tree Search for Efficient Information Seeking in Goal-Oriented Conversations

### v1

@@ Abstract
The ability to identify and acquire missing information is a critical component of effective decision making and problem solving. With the rise of conversational artificial intelligence (AI) systems, strategically formulating information-seeking questions becomes crucial and demands efficient methods to guide the search process. We introduce a novel approach to adaptive question-asking through a combination of Large Language Models (LLM) for generating questions that maximize information gain, Monte Carlo Tree Search (MCTS) for constructing and leveraging a decision tree across multiple samples, and a hierarchical feedback mechanism to learn from past interactions. We present two key innovations: (1) an adaptive MCTS algorithm that balances exploration and exploitation for efficient search over potential questions; and (2) a clustering-based feedback algorithm that leverages prior experience to guide future interactions. Each incoming sample is assigned to a cluster based on its semantic similarity with previously observed samples. Our UCT (Upper Confidence bound for Trees) formulation selects optimal questions by combining expected rewards, a function of information gain, with a cluster-specific bonus that decays with depth, to emphasize the importance of early-stage questions that have proven effective for narrowing the solution space in similar samples. Experiments across three domains, including medical diagnosis and troubleshooting, demonstrate that our method leads to an average of 12% improvement in success rates and a 10x reduction in the average number of LLM calls made per conversation for the search process, in comparison to the state of the art. [^1] https://github.com/harshita-chopra/misq-hf
@@ 1 Introduction
The field of conversational AI has been witnessing significant advancements in recent years, with a growing focus on developing systems capable of adaptive and efficient information-seeking strategies. As Large Language Models (LLMs) continue to push the boundaries of natural language understanding and generation, researchers are exploring novel approaches to enhance their decision-making capabilities in interactive contexts. One of the core challenges in a conversational system that employs multiple rounds of Q&A to arrive at an answer is to determine the next question to ask. A common solution to this challenge is to construct a contextual decision tree that allows the system to strategically plan which questions to pose and in what order, refining the process to ensure the correct answer or goal is achieved efficiently.
LLMs have emerged as powerful tools for planning using trees Zhao et al. (2024); Putta et al. (2024); Hu et al. (2024). Notably, the Tree of Thoughts (ToT) Yao et al. (2023) approach has demonstrated the potential of leveraging hierarchical structures to improve problem-solving and reasoning in language models. However, expanding the full tree can be computationally inefficient and expensive, especially in domains with large search spaces. One of the fundamental problems lies in designing systems capable of dynamically adapting their interaction strategy to converge on a specific solution efficiently. To address these challenges, Monte Carlo Tree Search (MCTS) has emerged as a promising technique for balancing exploration and exploitation in decision-making processes. MCTS has been successfully applied in various domains, from game-playing AI Silver et al. (2016, 2018) to robotics Dam et al. (2022), and is now being leveraged in natural language processing tasks.
Figure 1: Overview of MISQ-HF. The system combines Monte Carlo Tree Search with hierarchical feedback for efficient information seeking. Green nodes ( ) and red nodes ( ) represent $v^{A}$ and $v^{N}$ respectively, corresponding to possibilities after affirmative and negative answers to a question. The purple arrows indicate the information-seeking question loop (steps 4-8). Towards the end of decision-making, targeted questions are generated, denoted by the loop of dashed arrows (steps 9-6-7-8). Finally, when the user response $a_{t}$ (step 7) confirms the target detection, the conversation is terminated (step 10) and the feedback mechanism updates the reward bonus along the successful questioning trajectory in the decision tree, enabling the system to adapt its question selection strategy for similar future cases.
Building upon this foundation, our work introduces a novel framework that combines MCTS with a hierarchical feedback mechanism to achieve adaptive and efficient information-seeking. Inspired by prior work on uncertainty-aware planning, our approach extends these ideas with three key contributions:
@@
    MCTS on Decision Tree for Question Generation: Traditional tree-based approaches can struggle with the vast search spaces. By employing MCTS, our system efficiently constructs and explores promising branches of the decision tree without exhaustive expansion. We propose a variant of MCTS that optimizes question selection by considering both immediate and long-term information gain. We use the standard Upper Confidence Bound for Trees (UCT) and propose a modified variant that assigns higher weights to questions that have effectively narrowed the solution space in past interactions.
@@
    Cluster-Based Feedback Mechanism: We leverage similarity-based clustering based on text embeddings of the initial problem description given by the user. We allow the system to learn from historical successes by propagating cluster-specific bonus rewards through the decision tree after a successful interaction. This bonus reward decays with depth, recognizing that early-stage questions play a crucial role in narrowing down the solution space by targeting high-level information that can effectively eliminate large subsets of possibilities. These questions are often broadly applicable across multiple cases, making them valuable for generalizing strategies. In contrast, later-stage questions tend to be more specific to individual cases, focusing on finer details that depend on the unique context of each interaction.
@@
    Efficiency through reduced LLM calls: Our method minimizes computational overhead by strategically limiting LLM interactions during tree construction and simulation. We maintain a single decision tree for a given dataset which expands as more data points stream in. We also employ lightweight rollout policies limited by predefined depth during the simulation step of MCTS. The exploration-exploitation balance avoids exhaustive tree expansion and significantly reduces the number of LLM calls without compromising performance.
@@
We evaluate our approach across diverse conversational tasks. Results demonstrate that our system outperforms existing baselines in both task efficiency and relevance, particularly in scenarios requiring complex reasoning and hierarchical decision-making. Additionally, we also highlight the individual contributions of depth-aware MCTS and cluster-based feedback in enhancing the system’s performance.
@@ 2 Related Work
The integration of LLMs with tree search algorithms has proven effective for enhancing multi-step reasoning and decision-making. Guez et al. (2018) introduced MCTSnets, combining tree search with neural networks for tasks like Sokoban. While it focuses on discrete planning, our approach extends MCTS to hierarchical conversations, optimizing question selection with depth-aware strategies and semantic embeddings. Yu et al. (2023) used MCTS for dialogue planning with LLM-based simulations, but their approach differs in using open-loop search without feedback mechanisms and focuses only on persuasion tasks.
Hui and Tu (2024) introduced the Reflection on Search Trees (RoT) framework for generating state-specific guidelines to improve search performance. Unlike RoT, which focuses on reasoning efficiency, our approach maintains a single decision tree while incorporating cluster-specific rewards to guide question selection. Zhu et al. (2021) framed retrieval and answering as a partially observed Markov decision process to adaptively gather evidence, though it does not address learning from feedback in conversational settings. In the clinical domain, Li et al. (2024) proposed MEDIQ, an adaptive framework leveraging LLMs to identify missing information and ask follow-up questions. While MEDIQ emphasizes reliability in high-stakes contexts, our work generalizes by prioritizing early-stage questions to effectively narrow the solution space.
Xie et al. (2024) enhanced LLM reasoning through iterative preference learning inspired by AlphaZero, using MCTS to collect step-level preferences. This aligns with our approach of integrating cluster-specific success bonus rewards into MCTS. Zhang et al. (2024) explored autonomous tree search capabilities in LLMs, demonstrating self-guided reasoning, while our method emphasizes leveraging historical feedback through clustering mechanisms.
Our work draws inspiration from the Uncertainty of Thoughts (UoT) algorithm introduced by Hu et al. (2024), which enhances LLMs’ ability to actively seek information by asking effective questions. UoT combines a simulation approach with uncertainty-based rewards motivated by information gain, and a reward propagation scheme to select optimal questions, demonstrating significant improvements in task performance across various domains and baselines like Chain of Thoughts Wei et al. (2022) and Tree of Thoughts Yao et al. (2023). Building upon these principles, our approach extends the concept by integrating a depth-aware MCTS algorithm that optimizes question selection in hierarchical conversations, and a cluster-based feedback system that leverages historical successes to guide future interactions by adding a depth-aware bonus reward to prioritize early-stage questions. By integrating these components, our approach achieves a dynamic balance between efficiency and adaptability.
@@ 3 Methodology
In this section, we present our problem formulation, construction of the decision tree, reward function, our MCTS algorithm, and our hierarchical feedback approach.
@@ 3.1 Problem Formulation
The task is modeled as a sequential interaction between a Questioner (an LLM) and an Answerer (a human; here, simulated by an LLM). Let $\Omega$ be the space of possible target items, with an unknown target $\omega\in\Omega$. The interaction proceeds over turns $t=1,\dots,T$, where $T$ is the maximum allowed turns. At each turn, the system asks a question $q_{t}$ and receives an answer $a_{t}$, which is either binary (yes/no) or open-ended. The history $h_{t}=\{q_{1},a_{1},\dots,q_{t-1},a_{t-1}\}$ represents all previous interactions. At each step $i$, the current possibility set $\Omega_{i}$ is updated based on prior interactions. The updated set $\Omega_{i}$ contains all elements consistent with the history $h_{t}$. Questions in the first $\delta*T$ turns are selected strategically by traversing the decision tree of potential questions, and the remaining turns are reserved for making informed decisions about what the outcome is, based on $h_{i}$ and $\Omega_{i}$. Here $\delta$ denotes the proportion of turns reserved for asking information-seeking questions. The process ends when the Questioner identifies $\omega$ or reaches the maximum number of turns $T$.
Upon reaching terminal states in the decision tree (where $|\Omega_{t}|$ $\leq 2$), the Questioner LLM transitions from asking information-seeking questions to making targeted questions about specific outcomes. In the Closed Set scenario, where the target is one of the items within a predefined set $\Omega$, the tree construction is well-defined. However, this space is often unknown in real-world situations, leading to an Open Set scenario, where the models operate without prior knowledge of the outcomes. To address this, we follow the approach used in Hu et al. (2024), where we directly prompt the Question Generator to first define the initial possibility space $\Omega$ based on the problem description and then update it progressively based on the history of interactions $h_{t}$.
In practice, it is important to declare the set of all possible items $\Omega$ in the initial prompt for the Questioner LLM to avoid longer dialogues and failures due to wild guesses in the Closed Set scenario. Hence, to maintain logical consistency, we ensure that the Questioner LLM is explicitly informed about $\Omega$ once at the beginning of each interaction. This helps in constraining the items guessed in the targeting questions. To justify this modification, we report results with and without the $\Omega\text{-}aware$ prompt.
@@ 3.2 Decision Tree of Questions
The question generation process employs a hierarchical approach that combines LLMs with Monte Carlo Tree Search (MCTS) to efficiently construct a tree of potential questions represented by nodes. In the information-seeking phase, at each turn $t$, the system either traverses or expands the tree to find the most optimal questions to ask, with the goal of maximizing information gain. This tree is cached and re-used across all data-points belonging to a given dataset.
The LLM-based question generator takes two inputs: the current possibility set $\Omega_{t}$ and the ancestral context $\mathcal{A}_{\Omega t}$ which consists of the sequence of question-answer pairs corresponding to all the ancestor nodes of the current node that led to the formation of $\Omega_{t}$. This context prevents the generation of redundant or previously asked questions, thereby providing a clearer path of deduction. Let $v$ represent a node in the decision tree at turn $t$. With slight abuse of notation, we define $\Omega_{v}$ as the possibility set corresponding to node $v$. For each non-terminal node $v$, where $|\Omega_{v}|>2$, the LLM generates $m$ candidate questions which form its immediate child nodes:
@@
For each generated question, $LLM_{gen}$ performs a binary partitioning of the possibility set $\Omega_{v}$ into two disjoint subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$, corresponding to affirmative and negative responses respectively, such that $\Omega_{v}=\Omega_{v}^{A}\cup\Omega_{v}^{N}$. The LLM is prompted to generate questions that maximize information gain by creating balanced partitions where:
$|\Omega_{v}^{A}|-|\Omega_{v}^{N}|\rightarrow 0$ (2)
Each question spawns two child nodes $v^{A}$ and $v^{N}$ corresponding to affirmative and negative responses, storing $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ respectively. For example, if the possibility set is $\Omega_{v}=\{\text{flu},\text{pneumonia},\text{enteritis},\text{asthma}\}$, the question “Do you have difficulty breathing?” might partition the set as: $\Omega_{v}^{A}=\{\text{pneumonia},\text{asthma}\},\quad\Omega_{v}^{N}=\{\text{flu},\text{enteritis}\}$.
3.3 Uncertainty-Based Reward
To quantify the effectiveness of each question in reducing uncertainty, entropy-based metrics are employed Hu et al. (2024). Following the formulation of reward structure in UoT, let $p_{v}^{A}=|\Omega_{v}^{A}|/|\Omega_{v}|$ and $p_{v}^{N}=1-p_{v}^{A}$. The expected information gain at $v$ is:
@@ 3.3 Uncertainty-Based Reward
The reward function, given by $R_{IG}(v)$, achieves its maximum value when the subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ have equal probabilities, signifying the greatest reduction in uncertainty. It attains its minimum value when one of the subsets has a probability of zero, indicating no reduction in uncertainty. To normalize and sharpen rewards, a scaling parameter $\lambda>0$ is employed. Formally, $R_{IG}(v)$ is defined as:
$R_{IG}(v)=\frac{-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N}}{1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|}.$ (4)
The expected reward of asking a question, $R_{e}(v)$, is obtained by recursively adding the immediate reward $R_{IG}(v)$ (i.e., information gain at the current node) and the expected rewards of its child nodes. The total information gained until node $v$ of an interaction trajectory is denoted by the accumulated reward, $R_{a}(v)$, which is obtained by starting at the root and propagating down to the node $v$.
@@
where $v^{A}$ and $v^{N}$ are the child nodes of $v$ corresponding to affirmative and negative responses respectively. For each non-terminal response-specific node $v^{\alpha}$, $\alpha\in\{A,N\}$, the expected reward is computed as the average of the expected rewards of its child nodes, denoted by $\textit{children}(v^{\alpha})$. Formally:
@@
3.4 Monte Carlo Tree Search

### v2

@@ Abstract
Effective decision-making and problem-solving in conversational systems require the ability to identify and acquire missing information through targeted questioning. A key challenge lies in efficiently narrowing down a large space of possible outcomes by posing questions that minimize uncertainty. To address this, we introduce a novel framework that leverages Large Language Models (LLMs) to generate information-seeking questions, with Monte Carlo Tree Search (MCTS) to strategically select questions that maximize information gain, as a part of inference-time planning. Our primary contribution includes a hierarchical feedback mechanism that exploits past interaction patterns to guide future strategy. Specifically, each new problem is mapped to a cluster based on semantic similarity, and our UCT (Upper Confidence bound for Trees) formulation employs a cluster-specific bonus reward to prioritize successful question trajectories that have proven effective for similar problems in the past. Extensive empirical evaluation across medical diagnosis and technical troubleshooting domains shows that our method achieves an average of 12% improvement in success rates and about 10x reduction in the number of LLM calls made for planning per conversation, compared to the state of the art. An additional 8% gain in success rate is observed on average when we start with a constrained set of possibilities. Our results underscore the efficacy of feedback-aware MCTS in enhancing information-seeking in goal-oriented dialogues.  [^1] Our code is available here.
@@ 1 Introduction
Recent advances in conversational AI have transformed human-machine interactions, enabling more sophisticated goal-oriented dialogue systems. A fundamental requirement in these systems is the ability to identify and seek out missing information efficiently. When confronted with an initially under-specified problem, conversational agents must interact over multiple turns and strategically ask questions that reduce uncertainty while minimizing interaction overhead. This necessitates a principled approach to sequential decision-making wherein each question posed must maximize information gain and narrow down the solution search space. Large Language Models (LLMs) have shown strong capabilities in natural language understanding and generation, offering a promising basis for enhancing these information-seeking strategies. However, effectively applying LLMs to domains that involve complex problem-solving still requires overcoming challenges in efficient planning, exploration-exploitation trade-offs, and adaptation to domain-specific interaction patterns.
LLMs have emerged as powerful tools for planning using trees [22, 12, 7]. Notably, the Tree of Thoughts (ToT) [20] approach has demonstrated the potential of leveraging hierarchical structures to improve problem-solving and reasoning in language models. However, expanding the full tree can be computationally inefficient and expensive, especially in domains with large search spaces. A fundamental challenge lies in designing systems capable of dynamically adapting their interaction strategy to converge on a specific solution efficiently. To address these challenges, Monte Carlo Tree Search (MCTS) [2] has emerged as a promising technique for balancing exploration and exploitation in decision-making processes. It has been successfully applied in various domains, from game-playing AI [15, 16] to robotics [3], and is now being leveraged in natural language processing tasks.
Figure 1: Overview of MISQ-HF. In the MCTS block, green nodes ( ) and red nodes ( ) represent $v^{A}$ and $v^{N}$ respectively, corresponding to possibilities after affirmative and negative answers to a question. The purple arrows indicate the information-seeking question loop (steps 4-8). Towards the end of decision-making, targeted questions are generated, denoted by the loop of dashed arrows (steps 9-6-7-8). Finally, when the user response $a_{t}$ (step 7) confirms the target detection, the conversation is terminated (step 10) and the feedback mechanism updates the reward bonus along the successful questioning trajectory in the decision tree, enabling the system to adapt for similar future cases.
Building upon this foundation, our work introduces a novel framework that combines MCTS with a hierarchical feedback mechanism to achieve adaptive and efficient information-seeking. Inspired by prior work on uncertainty-aware planning, our approach presents three key contributions:
@@
    MCTS for Question Selection: LLMs can generate multiple valid questions to narrow search spaces, but identifying the optimal ones is crucial. Tree-based approaches are known for comprehensive planning. By employing MCTS, our system efficiently constructs and explores promising questions in the decision tree without exhaustive expansion, focusing on both immediate and long-term information gain. Our modified variant of the Upper Confidence Bound for Trees (UCT) assigns higher weights to questions that have effectively narrowed the solution space in past interactions.
@@
    Cluster-Based Feedback Mechanism: We cluster similar cases using text embeddings derived from the user’s initial problem description. After a successful interaction, the system propagates cluster-specific bonus rewards through the decision tree. This bonus reward decays with depth, emphasizing the importance of early-stage questions that effectively reduce the solution space by targeting broad, generalizable information.
@@
    Efficiency through reduced LLM calls: Our method minimizes computational overhead by strategically limiting LLM interactions during tree construction and simulation. We maintain a single decision tree for a given dataset, which expands as more data points stream in. We also employ lightweight rollout policies limited by predefined depth during the simulation step of MCTS. The exploration-exploitation balance avoids exhaustive tree expansion and significantly reduces the number of LLM calls without compromising performance.
@@
We evaluate our approach across diverse conversational tasks. Results demonstrate that our system outperforms existing baselines in both task success and efficiency in scenarios requiring complex reasoning and hierarchical decision-making. We also highlight the individual contributions of depth-aware MCTS and cluster-based feedback in enhancing the system’s performance.
@@ 2 Related Work
The integration of LLMs with tree search algorithms has proven effective for enhancing multi-step reasoning and decision-making. Guez et al.[5] introduced MCTSnets, combining tree search with neural networks for tasks like Sokoban. While it focuses on discrete planning, our approach extends MCTS to hierarchical conversations, optimizing question selection with depth-aware strategies and semantic embeddings. Yu et al.[21] applied MCTS for dialogue planning with LLM-based simulations, but their method differs by using open-loop search without feedback mechanisms and focusing solely on persuasion tasks. Hui et al.[8] proposed the Reflection on Search Trees (RoT) framework for generating state-specific guidelines to improve search performance. Unlike RoT, which prioritizes reasoning efficiency, our approach maintains a single decision tree while incorporating cluster-specific rewards to guide question selection. Zhu et al.[23] framed retrieval and answering as a partially observed Markov decision process to adaptively gather evidence but does not address learning from feedback in conversational settings. In the clinical domain, Li et al.[10] introduced MEDIQ, an adaptive framework leveraging LLMs to identify missing information and ask follow-up questions. While MEDIQ focuses on reliability in high-stakes contexts, our method generalizes by prioritizing early-stage questions to narrow the solution space effectively.
Our work draws inspiration from the Uncertainty of Thoughts (UoT) algorithm [7], which enhances LLMs’ ability to seek information through effective questioning. UoT uses uncertainty-based rewards driven by information gain and a reward propagation scheme to optimize question selection, improving performance across various domains and baselines like Chain of Thoughts [18] and Tree of Thoughts [20]. Extending these principles, we integrate a depth-aware MCTS algorithm for question selection in hierarchical conversations and a cluster-based feedback system that leverages historical successes. The depth-aware bonus reward prioritizes early-stage questions, balancing efficiency and adaptability.
@@ 3.1 Problem Formulation
The task is modeled as a sequential interaction between a Questioner (an LLM) and an Answerer (a human; here, simulated by an LLM). Let $\Omega$ be the space of possible target outcomes, with an unknown target $\omega\in\Omega$. The interaction proceeds over turns $t=1,\dots,T$, where $T$ is the maximum allowed turns. At each turn, the system asks a question $q_{t}$ and receives an answer $a_{t}$, which is either binary (yes/no) or open-ended. The history $h_{t}=\{q_{1},a_{1},\dots,q_{t-1},a_{t-1}\}$ represents all previous interactions. At each step $t$, the current possibility set $\Omega_{t}$ is updated based on prior interactions. The updated set $\Omega_{t}$ contains all elements consistent with the history $h_{t}$. Questions in the first $\delta*T$ turns, where $\delta\in(0,1)$, are selected strategically by traversing the decision tree of potential questions, and the remaining turns are reserved for making informed decisions about what the outcome (target) is, based on $h_{t}$ and $\Omega_{t}$. Here, $\delta$ denotes the proportion of turns reserved for asking information-seeking questions. The process ends when the Questioner identifies $\omega$ or reaches the maximum turns $T$.
Upon reaching terminal states in the decision tree (where $|\Omega_{t}|$ $\leq 2$), the Questioner LLM transitions from asking information-seeking questions to making targeted questions about specific outcomes. In the Closed Set scenario, where the target is one of the items within a predefined set $\Omega$, the tree construction is well-defined. However, this space is often unknown in real-world situations, leading to an Open Set scenario, where the models operate without prior knowledge of the outcomes. To address this, we follow the approach used in [7], where we directly prompt the Question Generator to first define the initial possibility space $\Omega$ based on the problem description and then update it progressively based on the history of interactions $h_{t}$. In practice, it is important to declare the set of all possible items $\Omega$ in the initial prompt for the Questioner LLM to avoid longer dialogues and failures due to arbitrary guesses in the Closed Set scenario. Hence, we ensure that the Questioner LLM is explicitly informed about $\Omega$ once at the beginning of each interaction. This helps fixing the items that appear in the targeting questions. To justify this modification, we report results with and without the $\Omega\text{-}aware$ prompt.
@@ 3.2 Decision Tree of Questions
The question generation process employs a hierarchical approach that combines LLMs with MCTS to efficiently construct a tree of potential questions represented by nodes. In the information-seeking phase, at each turn $t$, the system either traverses or expands the tree to find the most optimal questions to ask, with the goal of maximizing information gain. This tree is cached and reused across all data-points belonging to a given dataset.
The root node represents the initial conversational state in the decision tree, containing the complete predefined set of possibilities $\Omega$ in a dataset or specific domain. Over each of the question-answer turns, some of these possibilities are eliminated by asking information-seeking questions. Starting with a common, fixed superset of possibilities allows for efficiency, as the generated questions (child nodes) can be reused for future samples. This can lead to more general or broadly applicable questions in the beginning. However, for more relevant questions, we can initialize the root node as a constrained set of possibilities $\Omega_{c}\subseteq\Omega$, conditioned by a user’s problem description. Specifically, when a new problem arrives, the LLM is prompted to classify likely and unlikely outcomes in $\Omega$ based on the user’s problem. Upon determining $\Omega_{c}$, we search the existing first layer of tree for a root node whose possibility set $\Omega_{root}$ is a superset of the constrained set, i.e., $\Omega_{root}\supseteq\Omega_{c}$. If such a root node is found, we initiate the questioning process from that node, leveraging the existing question paths or generating more from there. Otherwise, a new root node is added to the tree, initialized with the constrained set $\Omega_{c}$. This can lead to a wider tree, as there are multiple possible root nodes, essentially representing different subsets of $\Omega$. We report results with both types of initializations.
The LLM-based question generator takes two inputs at turn $t$: the current possibility set $\Omega_{t}$ and the ancestral context $\mathcal{A}_{\Omega t}$ which consists of the sequence of question-answer pairs corresponding to all the ancestor nodes of the current node that led to the formation of $\Omega_{t}$. This context prevents the generation of redundant or previously asked questions, thereby providing a clearer path of deduction. Let $v$ represent a node in the decision tree at turn $t$. With slight abuse of notation, we define $\Omega_{v}$ as the possibility set corresponding to node $v$. For each non-terminal node $v$, where $|\Omega_{v}|>2$, the LLM generates $m$ candidate questions which form its immediate child nodes:
@@
For each generated question, $LLM_{gen}$ performs a binary partitioning of the possibility set $\Omega_{v}$ into two disjoint subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$, corresponding to affirmative and negative responses respectively, such that $\Omega_{v}=\Omega_{v}^{A}\cup\Omega_{v}^{N}$. The LLM is prompted to generate questions that maximize information gain by creating balanced partitions where: $|\Omega_{v}^{A}|-|\Omega_{v}^{N}|\rightarrow 0$.
Each question spawns two child nodes $v^{A}$ and $v^{N}$ corresponding to affirmative and negative responses, storing $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ respectively. For example, if the possibility set is $\Omega_{v}=\{\text{flu},\text{pneumonia},\text{enteritis},\text{asthma}\}$, the question “Do you have difficulty breathing?" will partition the set as: $\Omega_{v}^{A}=\{\text{pneumonia},\text{asthma}\},\quad\Omega_{v}^{N}=\{\text{flu},\text{enteritis}\}$.
3.3 Information Gain
To quantify the effectiveness of each question in reducing uncertainty, entropy-based metrics are employed ([7]). Following the formulation of reward structure in UoT, let $p_{v}^{A}=|\Omega_{v}^{A}|/|\Omega_{v}|$ and $p_{v}^{N}=1-p_{v}^{A}$. The expected information gain at $v$ is:
@@ 3.3 Information Gain
The reward function, given by $R_{IG}(v)$, achieves its maximum value when the subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ have equal probabilities, signifying the greatest reduction in uncertainty. It attains its minimum value when one of the subsets has a probability of zero, indicating no reduction in uncertainty. To normalize and sharpen rewards, a scaling parameter $\lambda>0$ is employed. Formally,
$R_{IG}(v)=\frac{-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N}}{1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|}$ (3)
The expected reward of asking a question, $R_{e}(v)$, is obtained by recursively adding the immediate reward $R_{IG}(v)$ (information gain at the current node) and the expected rewards of its child nodes. The total information gained until node $v$ of an interaction trajectory is denoted by the accumulated reward, $R_{a}(v)$, which is obtained by starting at the root and propagating down to the node $v$.
@@
where $v^{A}$ and $v^{N}$ are the child nodes of $v$ corresponding to affirmative and negative responses respectively. For each non-terminal response-specific node $v^{\alpha}$, $\alpha\in\{A,N\}$, the expected reward $R_{e}(v^{\alpha})$ is the average expected reward of its child nodes, denoted as $\textit{children}(v^{\alpha})$. Formally,
@@
3.4 Monte Carlo Tree Search (MCTS)